# STG-Mamba: Spatial-Temporal Graph Learning via Selective State Space Model

### v3

@@ 4.3 Results Evaluation and Comparison
A comprehensive evaluation between STG-Mamba and the proposed baseline methods is conducted, and we report the RMSE/MAE/MAPE performance results in Table 2. Our STG-Mamba consistently outperforms other benchmark methods on the three tasks, except for the MAPE criteria of PeMS04 (Flow). In PeMS04 (Flow) forecasting task, STAEformer shows slightly better MAPE results than STG-Mamba. It can also be easily observed that Transformer-based methods which integrate GNNs inside the architecture always gain superior performance than sole GNN-based methods. For instance, PDFormer, STAEformer, and MultiSPANS always show better performance than STGCN, STSGCN, and STG-NCDE. The reason is that although GNN is an effective approach of modeling STG tasks, Transformer-based methods have stronger ability to capture local and long-term dynamic dependencies. Within the GNN-based methods, STG-NCDE and DDGCRN stand out for their competitive performance. STG-NCDE models continuous-time dynamics within STG networks through neural controlled differential equations (NCDE), offering a more precise representation of spatiotemporal feature changes over time compared to other discrete GNN-based methods. Figure 4 presents a direct performance visualization comparison among STG-Mamba and other superior Transformer-based baselines (ASTGNN/PDFormer/STAEformer) on PeMS04, HZMetro, and KnowAir dataset, respectively. We visualize the predictions over 72 consecutive time steps using the average results of multiple 12-step forecast settings. For each dataset, four node individuals are randomly selected (i.e., 4 random traffic sensors of PeMS04, 4 random metro stations of HZMetro, and 4 random weather stations of KnowAir). The results indicate that STG-Mamba consistently yields the prediction curve closest to the ground truth waveform, featuring more accurate STG sequence peak values and fewer fluctuations.
Figure 4: STG forecasting results visualization and comparison among STG-Mamba and other superior Transformer-based baselines.
As a new contender poised to challenge Transformer architecture, Mamba introduces an innovative approach by replacing the foundational Attention mechanism with modern selective state space model (SSSM). SSSM diverges from Attention by utilizing a continuous-time model that captures the dynamic evolution of spatial-temporal dependencies more naturally and efficiently. Unlike Attention, which computes weights across all pairs of input and output positions and can be computationally intensive, Mamba’s selective state space model directly captures the temporal evolution of features, enabling it to scale more gracefully with the sequence length and complexity of spatial-temporal data. Furthermore, we facilitate STG-Mamba with Deep Kalman Filtering-based adaptive spatiotemporal graph evolution, and the elaborately designed Graph Selective State Space Block (GS3B), making it inherit the advantages of modern SSSM, and highly suitable for STG learning tasks.
Table 2: Performance Evaluation and Comparison with Benchmark Baselines on PeMS04, HZMetro, KnowAir Datasets.
@@
RMSE MAE MAPE RMSE MAE MAPE RMSE MAE MAPE
STGCN 35.55 22.70 14.59 34.85 21.33 13.47 20.64 15.27 12.92
STSGCN 33.65 21.19 13.90 33.24 20.79 13.06 20.25 15.22 12.69
STG-NCDE 31.09 19.21 12.76 32.91 20.75 12.88 20.08 14.93 11.87
DDGCRN 30.51 18.45 12.19 31.69 19.52 12.45 19.63 12.45 11.84
ASTGCN 35.22 22.93 16.56 34.36 21.12 13.50 20.37 15.07 12.11
ASTGNN 31.16 19.26 12.65 32.75 20.63 12.47 19.96 12.38 11.23
PDFormer 29.97 18.32 12.10 30.18 19.13 11.92 18.65 12.21 11.06
STAEformer 30.18 18.22 11.98 29.94 18.85 12.03 18.49 11.93 10.75
MultiSPANS 30.46 19.07 13.29 30.31 18.97 11.85 18.57 12.24 10.88
STG-Mamba 29.53 18.09 12.12 29.26 18.38 11.67 18.05 11.73 10.66
@@ 4.4 Robustness Analysis
Table 3: Robustness analysis on PeMS04. We compare the forecasting performance among STG-Mamba and 4 superior baseline methods during rush/non rush hours, weekend/non weekend settings.
Models Rush Hours Non Rush Hours Weekend Non Weekend
RMSE MAE MAPE RMSE MAE MAPE RMSE MAE MAPE RMSE MAE MAPE
ASTGNN 33.57 21.05 13.81 29.76 18.22 12.49 30.66 19.08 13.35 31.82 19.53 12.28
PDFormer 31.69 19.54 12.38 28.95 18.16 11.97 29.87 18.26 12.04 30.39 18.68 12.23
STAEformer 31.36 19.10 12.22 29.47 17.98 11.93 29.83 18.19 11.96 30.26 18.42 12.27
MultiSPAN 30.95 19.28 13.31 30.04 18.37 12.54 30.13 18.42 12.73 30.52 19.31 13.34
STG-Mamba 29.58 18.13 12.15 29.23 18.04 11.91 29.31 18.05 11.92 29.57 18.15 12.16
@@
STG data exhibit notable periodicity and diversity. For urban mobility/traffic data, there are distinct differences in traffic characteristics between the peak travel periods in the morning and evening and the none-peak periods. Moreover, the traffic patterns over the 24 hours of weekdays and those of weekends also differ substantially. In light of the variations in STG characteristics due to external environmental changes, it is imperative to assess whether deep learning models can accurately and effectively model the spatial-temporal dependencies under various external conditions.
As such, we establish four distinct external scenarios: (a) Traffic rush hours: Monday to Friday, 8:00-11:00 AM and 16:00-19:00 PM; (b) Non-rush hours: encompassing the remaining time periods from Monday to Friday; (c) Weekend: including all hours on Saturday and Sunday; and (d) Non-weekend: including all time periods from Monday to Friday. Extensive experiments were conducted on the PeMS04 dataset. Table 3 presents the STG forecasting results corresponding to the four traffic scenarios. We compare the prediction performance of STG-Mamba with the other four superior baselines, and report the RMSE/MAE/MAPE metrics.
An ideal deep learning-based STG system should maintain robustness to external disturbances (changes of external environment or events). This is specifically reflected in the fact that when facing different scenarios, the model performance should maintain as little difference as possible, rather than being good at some scenarios but being bad at other occasions.
Among the compared methods, ASTGNN, PDFormer, and STAEformer exhibit relatively large differences in the prediction results between the Rush hour and Non-rush Hour scenarios. In the Weekend/Non Weekend scenario, although the performance differences of each model are not as significant as in the Rush hour/Non-Rush hour scenario, we observe that STG-Mamba still shows the smallest performance difference (the maximum robustness) and achieves the best RMSE/MAE/MAPE metrics. In summary, STG-Mamba proves the best robustness and outperforms existing baseline methods even in face of complex and diversity traffic scenes.
@@ 4.5 Ablation Study
To investigate the effectiveness of each model component within STG-Mamba, we further design three kinds of model variants, and evaluate their forecasting performance on PeMS04, HZMetro, and KnowAir dataset. The detailed description of these model variants are introduced in the following:
  1. 1.
    STG-Mamba: The full STG-Mamba model without any modification.
  2. 2.
    STG-Mamba w/o KF-Upgrading: We replace the KF-Upgrading module in KFGN with the simple sum average operation to fuse the three temporal branches.
  3. 3.
    STG-Mamba w/o Dynamic Filter Linear: We replace the proposed Dynamic Filter Linear in KFGN with the basic static graph convolution.
  4. 4.
    STG-Mamba w/o SSSM: The State Space Selection Mechansim (SSSM) is removed from the ST-S3M module.
  5. 5.
    STG-Mamba w/o ST-S3M: The whole ST-S3M module is removed from GS3B Encoder.
@@ 5 Conclusions
In this work, we for the first time introduce deep learning-based selective state space models (SSSM) for spatial-temporal graph learning tasks. We propose STG-Mamba that leverages modern SSSM for accurate and efficient STG forecasting. STG-Mamba is formulated following the stacked Residual Encoder architecture, with the ST-S3M module facilitating input-dependent graph evolution and node feature selection, successfully integrates STG network with SSSMs. Through the Kalman Filtering Graph Neural Networks (KFGN), the learned STG embeddings achieve a smooth optimization & upgrading based on statistical learning, aligning with the whole STG selective state space modeling process. Compared with Attention-based methods, STG-Mamba is able to achieve linear-time complexity and substantial reduction in FLOPs and test inference time. We conduct extensive empirical studies on three open-sourced STG datasets from urban traffic, metro In/Out flow, and weather records. Experimental results demonstrate the consistent superiority of STG-Mamba over other benchmark methods. We believe that STG-Mamba presents a brand new promising approach to general STG learning fields, offering competitive model performance and input-dependent context-aware feature selection under an affordable computational cost.

### v4

@@ 4.4 Results Evaluation and Comparison
A comprehensive evaluation between STG-Mamba and the proposed baseline methods is conducted, and we report the RMSE/MAE/MAPE/$std_{\text{MAE}}$ performance results in Table 2. STG-Mamba consistently outperforms other benchmark methods, except for the MAPE criteria of PeMS04 (Flow). It can also be easily observed that Transformer-based methods which integrate GNNs inside the architecture always gain superior performance than sole GNN-based methods. For instance, PDFormer, STAEformer, and MultiSPANS always show better performance than STGCN, STSGCN, and STG-NCDE. The reason is that although GNN is an effective approach of modeling STG tasks, Transformer-based methods have stronger ability to capture local and long-term dynamic dependencies. Within the GNN-based methods, STG-NCDE and DDGCRN stand out for their competitive performance. STG-NCDE models continuous-time dynamics through neural controlled differential equations (NCDE), offering a more precise representation of spatiotemporal feature changes over time compared to other discrete GNN-based methods.
As a new contender poised to challenge Transformer architecture, the selective state space mechanism of Mamba diverges from Attention by utilizing a continuous-time model that captures the dynamic evolution of spatial-temporal dependencies more naturally and efficiently. Unlike Attention, SSSM directly captures the temporal evolution of features, enabling it to scale more gracefully with the sequence length and complexity of STG data. Furthermore, we facilitate STG-Mamba with Kalman Filtering-based adaptive STG evolution, the elaborately designed graph selective scan algorithm, and the Feed-Forward connection, making it inherit the advantages of modern SSSM, and highly suitable for STG learning.
Table 2: Performance Eval and Comparison with Baselines on PeMS04/HZMetro/KnowAir Dataset.
@@
RMSE MAE MAPE $std_{\text{MAE}}$ RMSE MAE MAPE $std_{\text{MAE}}$ RMSE MAE MAPE $std_{\text{MAE}}$
STGCN 35.55 22.70 14.59 1.2068 34.85 21.33 13.47 1.1835 11.46 8.37 11.26 0.4358
STSGCN 33.65 21.19 13.90 1.1820 33.24 20.79 13.06 1.1249 11.22 8.15 10.89 0.4134
STG-NCDE 31.09 19.21 12.76 1.0736 32.91 20.75 12.88 1.1097 10.85 7.93 10.47 0.3922
DDGCRN 30.51 18.45 12.19 1.0365 31.69 19.52 12.45 1.0586 10.43 7.84 10.38 0.3794
ASTGCN 35.22 22.93 16.56 1.1853 34.36 21.12 13.50 1.1258 10.27 7.57 10.51 0.3745
ASTGNN 31.16 19.26 12.65 1.0583 32.75 20.63 12.47 1.0784 9.68 7.35 10.23 0.3569
PDFormer 29.97 18.32 12.10 1.0217 30.18 19.13 11.92 1.0205 9.46 7.12 10.06 0.3375
STAEformer 30.18 18.22 11.98 0.9678 29.94 18.85 12.03 0.9728 8.69 6.93 9.89 0.3083
MultiSPANS 30.46 19.07 13.29 0.9851 30.31 18.97 11.85 0.9931 8.57 6.84 10.05 0.3149
STG-Mamba 29.53 18.09 12.11 0.9218 29.23 18.26 11.59 0.9271 8.05 6.37 9.64 0.2648
@@ 4.5 Robustness Analysis
Table 3: Robustness analysis. We choose rush/non-rush, weekend/non-weekend traffic scenarios.
Models Rush Hours Non Rush Hours
RMSE MAE MAPE RMSE MAE MAPE
ASTGNN 33.57 21.05 13.81 29.76 18.22 12.49
PDFormer 31.69 19.54 12.38 28.95 18.16 11.97
STAEformer 31.36 19.10 12.22 29.47 17.98 11.93
MultiSPAN 30.95 19.28 13.31 30.04 18.37 12.54
STG-Mamba 29.58 18.13 12.15 29.23 18.04 11.91
Models Weekend Non Weekend
RMSE MAE MAPE RMSE MAE MAPE
ASTGNN 30.66 19.08 13.35 31.82 19.53 12.28
PDFormer 29.87 18.26 12.04 30.39 18.68 12.23
STAEformer 29.83 18.19 11.98 30.26 18.42 12.27
MultiSPAN 30.13 18.42 12.73 30.52 19.31 13.34
STG-Mamba 29.31 18.05 11.92 29.57 18.15 12.16
@@
STG data exhibit notable periodicity and diversity, with urban mobility/traffic data showing clear differences between peak travel times in the morning/evening and non-peak periods, as well as between weekdays and weekends. Given these variations due to external environmental changes, it is significant to determine whether deep learning models can effectively model the spatiotemporal dependencies under different conditions. As such, we establish four distinct external scenarios: (a) rush hours on weekdays from 8:00-11:00 AM and 4:00-7:00 PM; (b) non-rush hours during weekdays; (c) weekends (all hours); and (d) non-weekend (all hours). Extensive experiments were conducted on the PeMS04 dataset. Table 3 presents the forecasting results of the four traffic scenarios.
An ideal deep learning-based STG system should be robust to external disturbances, maintaining consistent performance across different scenarios. Compared methods like ASTGNN, PDFormer, and STAEformer show significant performance differences between Rush hour and Non-rush Hour scenarios. Even in the Weekend/Non-Weekend scenario, STG-Mamba demonstrates the smallest performance difference, showing maximum robustness and achieving the best RMSE/MAE/MAPE metrics. In summary, STG-Mamba exhibits superior robustness and outperforms existing baselines in diverse traffic conditions.
Statistical-based evaluation in spatiotemporal forecasting is indispensible as they provide a quantifiable measure of how well a model captures and predicts complex data dynamics over time and space. Specifically, $R^{2}$ and $\Delta_{\text{VAR}}$ help in assessing the accuracy and robustness of models under varying conditions, ensuring the reliability of predictions and effectively inform decision-making processes. The Statistical evaluation results are presented in Table 4. Here, STG-Mamba demonstrates superior performance across datasets, achieving the highest $R^{2}$ values and the lowest $\Delta_{\text{VAR}}$ scores, indicating the advancement of accuracy and efficiency in handling spatiotemporal dependencies.
Table 4: Statistical evaluation results.
Model PeMS04 (Flow) HZMetro KnowAir
$\Delta_{\text{VAR}}$ $R^{2}$ $\Delta_{\text{VAR}}$ $R^{2}$ $\Delta_{\text{VAR}}$ $R^{2}$
STGCN 0.23145 0.77069 0.25769 0.74385 0.49851 0.50487
STSGCN 0.20538 0.79456 0.21474 0.78862 0.47536 0.52831
STG-NCDE 0.18275 0.81973 0.19827 0.80492 0.46173 0.54125
DDGCRN 0.13291 0.87355 0.14796 0.85511 0.43185 0.56506
ASTGCN 0.17289 0.82698 0.18131 0.82134 0.42729 0.58461
ASTGNN 0.14863 0.84850 0.15765 0.84627 0.39726 0.60563
PDFormer 0.11849 0.88190 0.13259 0.87095 0.36182 0.64719
STAEformer 0.09631 0.90632 0.10383 0.89738 0.33215 0.66085
MultiSPANS 0.06218 0.93804 0.07596 0.92561 0.32569 0.67122
STG-Mamba 0.04387 0.95489 0.05148 0.94752 0.30182 0.69537
@@ 4.6 Ablation Study
To investigate the effectiveness of each model component within STG-Mamba, we further design five kinds of model variants, and evaluate their forecasting performance on PeMS04/HZMetro/KnowAir dataset: (I) STG-Mamba: The full STG-Mamba model without any modification. (II) STG-Mamba w/o KF-Upgrading: We replace the KF-Upgrading module in KFGN with the simple sum average operation to fuse the three temporal branches. (III) STG-Mamba w/o Dynamic Filter Linear: We replace the proposed Dynamic Filter Linear in KFGN with the basic static graph convolution. (IV) STG-Mamba w/o GSSSM: The Graph State Space Selection Mechansim (GSSSM) is replaced by the basic SSSM Gu and Dao (2023). (V) STG-Mamba w/o ST-S3M: The whole ST-S3M module is removed from GS3B Encoder.
@@ 5 Conclusions
In this work, we for the first time introduce deep learning-based selective state space models (SSSM) for spatial-temporal graph learning tasks. We propose STG-Mamba that leverages modern SSSM for accurate and efficient STG forecasting. In STG-Mamba, the ST-S3M module facilitates input-dependent graph evolution and feature selection, successfully integrates STG network with SSSMs. Through the Kalman Filtering Graph Neural Networks (KFGN), the learned STG embeddings achieve a smooth optimization & upgrading based on statistical learning, aligning with the whole STG selective state space modeling process. Compared with Attention-based methods, STG-Mamba achieves linear-time complexity and substantial reduction in FLOPs and inference time. Extensive empirical studies are conducted on three open-sourced STG datasets, demonstrating the consistent superiority of STG-Mamba over other benchmark methods. We believe that STG-Mamba presents a brand new promising approach to general STG learning fields, offering competitive model performance and input-dependent contextual feature selection under an affordable computational cost.
Appendix A ST-S3M Graph Selective Scan Algorithm & Parameter Calculation and Upgrade Algorithm
Algorithm 1 Graph Selective Scan Algorithm
Input: $u,\Delta^{*},\mathbb{A},\mathbb{B},\mathbb{C},\mathbb{D}$;
Output: $z$;
1:  Obtain the dim of $u$ as $(b,l,d_{\text{in}})$ the second dim of $\mathbb{A}$ as $n$;
[…]
6:  Discretize continuous parameters $\mathbb{A}$ and $\mathbb{B}$:
7:  $\text{delta}\mathbb{A}\leftarrow\exp(\text{einsum}(\Delta^{{}^{\prime}},\mathbb{A}))$, where $\text{delta}\mathbb{A}\in\mathbb{R}^{b\times l\times d_{\text{in}}\times n}$;
8:  $\text{delta}\mathbb{B}_{u}\leftarrow\text{einsum}(\Delta^{{}^{\prime}},\mathbb{B},u)$, where $\text{delta}\mathbb{B}_{u}\in\mathbb{R}^{b\times l\times d_{\text{in}}\times n}$;
9:  Initialize State $x$ as zeroes with dimension $\mathbb{R}^{b\times d_{\text{in}}\times n}$;
10:  for $i$ from $1$ to $l$ do
11:     $x\leftarrow\text{delta}\mathbb{A}[:,i]\times x+\text{delta}\mathbb{B}_{u}[:,i]$;
12:     $z\leftarrow\text{einsum}(x,\mathbb{C}[:,i,:])$;
13:     Append $z$ to the output list $z_{s}$;
14:  end for
15:  $z\leftarrow\text{stack}(z_{s})$, where $z\in\mathbb{R}^{b\times l\times d_{\text{in}}}$;
16:  Adding the direct gain $\mathbb{D}$:
17:  $z\leftarrow z+u\times\mathbb{D}$
18:  Return $z$.
Algorithm 2 State Space Selection Mechanism Parameter Calculation & Upgrade Algorithm.
Input: $x\in\mathbb{R}^{b\times l\times d_{\text{model}}}$, learnable model params $\mathbb{A}_{log}\in\mathbb{R}^{d_{\text{in}}\times n}$ and $\mathbb{D}\in\mathbb{R}^{d_{\text{in}}}$;
Output: $y\in\mathbb{R}^{b\times l\times d_{\text{inner}}}$;
1:  Compute input-independent params $\mathbb{A}\in\mathbb{R}^{d_{\text{in}}\times n}$ and $\mathbb{D}\in\mathbb{R}^{d_{\text{in}}}$:
2:  $\mathbb{A}=-\exp(\mathbb{A}_{log})$, $\mathbb{D}=\mathbb{D}$;
3:  Process input $x$ with Linear Layer $Input\_proj$:
5:  Dividing $x_{\text{dbl}}$ in order to get the input-dependent params $\Delta,\mathbb{B},\mathbb{C}$, where $\Delta\in\mathbb{R}^{b\times l\times dt_{\text{rank}}}$, $\mathbb{B},\mathbb{C}\in\mathbb{R}^{b\times l\times n}$;
8:  Employ the GraphSelectiveScan algorithm to proceed with $x,\Delta^{*},\mathbb{A},\mathbb{B},\mathbb{C},\mathbb{D}$:
9:  $y=\text{GraphSelectiveScan}(x,\Delta^{*},\mathbb{A},\mathbb{B},\mathbb{C},\mathbb{D})$;
10:  Return the Output $y$.
Appendix B Computational Efficiency Analysis and Comparison
B.1 Theoretical Computational Efficiency Analysis
B.2 Experimental Computational Efficiency Analysis
STG-Mamba offers substantial improvement in computational efficiency in addition to performance gain. In order to quantatively evaluate computational efficiency, we select the widely recognized Floating Point Operations (FLOPs) and Inference Time as criteria. For both of the two criteria, a smaller value indicates better model performance. STG-Mamba is compare with other Transformer-based benchmark methods: ASTGNN, PDFormer, and STAEformer, on the same testing set.
Table 5 showcases the inference time comparison on PeMS04 and KnowAir testing set, we report the average inference time of every 100 batch with Sequence_Length=12, Batch_Size=48. It can be seen that ASTGNN is relatively redundancy, with the lowest inference time among the compared Transformer-based methods. STG-Mamba gains 18.09% and 20.54% performance increase in inference time than the 2nd best performance model (STAEformer).
We further compare the computational FLOPs increment with accordance to the scale complexity of STG network. Specifically, we use the number of node individuals to represent the scale of STG network. In the experiments, we select different number of nodes scales: $[50,100,150,200,250,300]$ for PeMS04 and $[30,60,90,120,150,180]$ for KnowAir, respectively. STG-Mamba is compared with the benchmark STAEformer in every settings. Figure 4 visualizes the FLOPs evaluation under different STG scale complexity. We can easily find that STG-Mamba enjoys a linear increment in FLOPs, while STAEformer shows quadratic increment in FLOPs. At the initial small-scale setting, the FLOPs for STG-Mamba/STAEformer are quite similar. However, with the increasing of network complexity and scale, the difference becomes more significant. At the 300 node setting of PeMS04, the FLOPs for STG-Mamba and STAEformer is 60.85G and 134.92G (2.22 times more), respectively. At the 180 node setting of KnowAir, the FLOPs for STG-Mamba and STAEformer is 34.03G and 102.79G (3.54 times more), respectively.
Benefiting from the linear computational increment brought by modern selective state space models, STG-Mamba and alike models demonstrate stronger capability and computational efficiency for modeling large-scale STG data, which holds promise for addressing the computational overhead in future large-scale STG learning.
Table 5: Test inference time comparison with benchmark Transformer-based methods.
Model Inference Time on PeMS04 Inference Time on KnowAir
ASTGNN 3.285 s 3.018 s
PDFormer 2.431 s 2.357 s
STAEformer 2.106 s 1.933 s
STG-Mamba 1.547 s 1.481 s
performance $\uparrow$ +26.54% +23.38 %
Figure 4: (Computational Efficiency Eval) FLOPs comparison between STG-Mamba and STAEformer on PeMS04/KnowAir dataset under different STG node number settings.
Appendix C Evaluation Metrics Calculation Method
We employ RMSE/MAE/MAPE/$std_{\text{MAE}}$ as the deterministic evaluation metrics, all the indicators are smaller the better; $\rm{\Delta_{\text{VAR}}}$ (Change in Variance Accounted For) and $\rm{R^{2}}$ (Coefficient of Determination) as statistic-based evaluation metric, where $\rm{\Delta_{\text{VAR}}}$ is the smaller the better, $\rm{R^{2}}$ is the larger the better. The reported results in Experiment section are the average value of 10 runs.
Here, $std_{\text{MAE}}$ refers to the standard deviation of multiple MAE values in the testing stage. $std_{\text{MAE}}$ provides a measure of consistency for MAE results across multiple evaluation sets. A smaller $std_{\text{MAE}}$ indicates the model performs more stably under different conditions, with less variation in prediction errors. The calculation formula for $std_{\text{MAE}}$ is as follows: Step1-Calculate the average MAE $\mu$: $\mu=\frac{1}{n}\sum_{i=1}^{n}\text{MAE}_{i}$. Here, $n$ denotes the number of batches in the testing dataset, with the MAE value computed for each batch. $\text{MAE}_{i}$ is the MAE result of the $i$-th batch. Step2-Calculate the standard deviation: $\text{std}_{\text{MAE}}=\sqrt{\frac{\sum_{i=1}^{n}(\text{MAE}_{i}-\mu)^{2}}{n}}$.
$R^{2}$ refers to the Coefficient of Determination is a statistical index used for measuring the proportion of variance in the dependent variable that is predictable from the independent variables. $R^{2}$ provides a gauge of the effectiveness of the model in explaining the variations in the observed data. A higher $R^{2}$ value indicates the model has a higher predictive accuracy and explains a substantial proportion of the variance in the dependent variable. $R^{2}$ is formulated as follows: $R^{2}=1-\frac{\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}{\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}}$ where $y_{i}$ denotes the ground-truth values, $\hat{y}_{i}$ denotes the model predictions, $\bar{y}$ is the mean of the ground-truth values. $\sum(y_{i}-\hat{y}_{i})^{2}$ indicates the unexplained variance by the model, while the denominator, $\sum(y_{i}-\bar{y})^{2}$, is the total variance in the observed data.
Change in Variance Accounted For ($\Delta_{VAR}$) quantifies the improvement or regression in how much variance the model can account for compared to a previous models or baselines. This metric is particularly valuable in iterative model optimization where understanding the variance captured by model updates is crucial. $\Delta_{VAR}$ is calculates as follows: $\Delta_{VAR}=\frac{\text{Var}(y-\hat{y})}{\text{Var}(y)}$. $\text{Var}(y-\hat{y})$ computes the variance of residuals, and $\text{Var}(y)$ is the variance of the ground-truth data. The difference between the variance ratio of the current and previous models provides the change in variance accounted for, reflecting the model’s improvement in explaining the data variability.
Appendix D Formulation of Selective State Space Models for STG Learning
Appendix E Discussion of Limitations
STG-Mamba for spatial-temporal graph learning demonstrates significant advantages, such as superior forecasting accuracy, model robustness, computational efficiency, and the ability to effectively handle datasets with noise and uncertainty. Despite these promising results, there are some limitations to consider:
  • Although STG-Mamba has been extensively evaluated on three diverse open-sourced datasets (Weather/Traffic/Urban Metro), these datasets may not fully representing the wide range of possible spatiotemporal scenarios, potentially limiting the generalizability of the model. Future work should aim to test the model on a broader range of datasets from different geographical regions and temporal scales.
  • The model interpretability is another concern worthy of further investigation. Same as other deep learning-based methods, the black-box nature of Mamba architecture makes it not easy to interpret how specific predictions are made. Enhancing the interpretability deep learning-based SSSMs could be beneficial, particularly in applications where understanding the decision-making process is crucial.
  • While the novel integration of Kalman Filtering statistic learning-based optimization with dynamic graph neural networks helps in managing noise and uncertainty, the model’s performance can still be affected by the quality and completeness of input dataset. Developing robust data pre-processing techniques and improving the model’s ability to directly handle incomplete or noisy data would further enhance its reliability.
We believe by addressing the aforementioned limitations, the future iterations of STG-Mamba can achieve even greater accuracy, robustness, and applicability across a broader range of spatial-temporal graph learning tasks.
Appendix F Illustration of the STG feature selection procedure
Figure 5 is the illustration of the spatial-temporal state space selection mechanism that facilitates input-dependent feature selection and focusing.
Appendix G Related Works
Owning to the strong ability of capturing dynamic features across both temporal and spatial dimension, attention-based methods, such as Spatial-Temporal Transformer and its variants have recently been broadly implemented in spatial-temporal graph (STG) learning tasks. In the domain of STG traffic forecasting, seminal works have emerged that leverage these methods to enhance prediction accuracy. ASTGCN Guo et al. [2019] introduces an innovative approach combining spatial-attention & temporal attention mechanisms with graph neural networks. ADCT-Net Kong et al. [2024] presents to incorporate a dual-graphic cross-fusion Transformer for dynamic traffic forecasting. Huo et al. [2023] integrates hierarchical graph convolutional networks with Transformer networks to capture sophisticated STG relationships among traffic data. For weather forecasting, Liang et al. [2023] exemplifies how Transformer-based models can be utilized for accurate air quality forecasting across vast geographical areas. Chen et al. [2023] further demonstrates the power of GNNs in predicting air quality, emphasizing the importance of group-aware mechanisms in enhancing STG forecasting performance. In the realm of Social Recommendation, Zhang et al. [2023] introduces an Adaptive Disentangled Transformer (ADT) framework that optimizes the disentanglement of attention heads in different layers of Transformer architecture. Wei et al. [2023] explores the improved efficiency of multimedia recommendation by introducing modal-specific embeddings and a lightweight self-attention.
Although Transformer-based methods demonstrate notable improvements in STG relationship learning Chen et al. [2022], Cong et al. [2021], their modeling ability on large-scale STG networks and Long-Term Sequential Forecasting (LTSF) tasks is greatly hindered by the quadratic computational complexity $O(n^{2})$ provided by its core component–attention mechanism Zhou et al. [2021, 2022], Rampášek et al. [2022]. Moreover, attention mechanisms usually approximate the full-length historical data, or encode all the context features, which may not be necessary for modeling the long-term dependencies of STG learning tasks. As such, the research world has been actively looking for counterparts of Attention-based models.